# QKD security proofs for decoy-state BB84: protocol variations, proof techniques, gaps and limitations

### v1

@@ 5.4.6 Error verification and the correctness condition
As previously discussed in Section 4.3, care is necessary when interpreting the guarantees of error verification. In particular, when considering error verification based on 2-universal hashing, the hashing procedure only offers the guarantee that conditioned on the pre-amplification strings being different, then the probability that their hashes match is small. We stress that this is not the same as a statement that conditioned on the hashes matching, the probability of the pre-amplification strings differing is small —these two probabilities are a priori entirely unrelated, but conflating them is a common mistake. It is in fact possible to construct explicit counterexamples to the latter statement, i.e. attacks by Eve such that
[…]
Hence the pre-amplification strings may differ with high probability even conditioned on the hashes matching.
Fortunately, the correctness condition in the composable security definition does not require a bound on the latter probability; it only requires that the probability of the pre-amplification strings differing and the hashes matching is small. As noted in Section 4.3 (Eq. 10), a bound on the former probability is indeed sufficient to show that the correctness definition is satisfied. The framework of composable security then ensures that we have the desired composable properties of the protocol, despite the fact that even conditioned on the hashes matching, Alice and Bob’s keys may differ with high probability.
@@ 5.5 Variable-length protocols
For security proofs of variable-length protocols see Refs. [30, 11] for an approach involving the use of the postselection technique [48], see [57, Supplementary Note A], [58, Appendix B, F] for a security proof in the EUR framework, and [59, Chapter 3] [60] for the phase error correction based methods.
@@ 5.6.1 Resources used in QKD
While the RNGs and the authenticated classical channels can in principle both be constructed with information-theoretic security, for some QKD implementations one might instead wish to use constructions that only have computational security. There exist works considering this possibility and its connection to a concept known as everlasting security, but that will not be a focus of this work.
@@ 5.6.2 Variable-length of pre-amplification string
In theoretical analysis, sifting is typically modeled as setting the pre-amplification string register to a 00 or $\bot$ for each round that is discarded, which leads to the pre-amplification string register having a fixed length. Privacy amplification is then implemented by two-universal hashing on this register (of fixed-length) to some $l$ bits. Note that if the discarded rounds are set to $\bot$, then the hashing procedure is no longer applied to binary strings.
@@
In practice, sifting is often implemented by physically discarding the round, which leads to the pre-amplification string register having a variable number of bits. Then, privacy amplification is done by first looking at the number of bits in the pre-amplification string, and then implementing 2-universal hashing from those many bits to $l$ bits. This approach is preferred since it allows users to apply hashing on smaller, binary strings.
This difference constitutes a clear gap between the theoretical model and the practical implementation. The gap is further exacerbated by the fact that the implemented protocol does not represent a valid two-universal hashing procedure on the pre-amplification string register [^35] For instance, consider two all-zero strings of different lengths. If a linear 2-universal hash is applied on these strings, the probability of collision is $1$, even though the input strings are different.. This issue was first pointed out and addressed in Ref. [30, Section V], where a variant of the Leftover Hashing Lemma that is applicable to the variable-length pre-amplification strings case was proven. In particular, it was shown that as long as the positions of the discarded rounds are publicly announced, the theoretical analysis of the following three scenarios are equivalent:
@@
Note that the above problem does not arise in EUR and phase error correction based proof techniques. This is because in these frameworks, the pre-amplification string is only generated from detected signals. If the pre-amplification string is of variable length, then one can separately apply the Leftover Hashing Lemma for each possible length, and obtain a security proof similar to the one for variable-length protocols [58, Remark 17]. Alternatively, some protocols choose a random subset of fixed length for the pre-amplification string on which privacy amplification is done.
@@ 6.1 Proofs based on entropic uncertainty relations
Additionally, this work requires the source preparation to be perfect. However, there are several follow up works on the topic of source imperfections, that provide methods to obtain a suitable bound on the phase error rate, even in the presence of source imperfections [81, 82, 83, 84]. It also considers active basis choice by Bob, and assumes that the probability of detection in Bob’s setup is independent of basis. This assumption is a core issue for all EUR and phase error correction (see Section 6.2) based proofs for standard QKD, and was removed recently in Ref. [58].
6.1.2 Rusca et al., 2018 [65]
@@ 6.2 Phase error correction approach
Refs. [85, 86] showed that if there exists some constant $N_{\mathrm{const}}^{\mathrm{err}}\in\mathbb{N}$ that bounds the number of phase errors with high probability, then hashing [^38] There are some restrictions in the hash families considered in this proof; we shall discuss this point shortly. the string $Z_{A}^{m}$ in the state $\rho_{Z_{A}^{m}\mathbf{E}}^{(Z)}$ to a length determined by $N_{\mathrm{const}}^{\mathrm{err}}$ results in the output being nearly uniform and decoupled from $\mathbf{E}$, roughly analogous to the Leftover Hashing Lemma. [^39] There are a number of similarities between this approach and EUR-based approaches when constructing a QKD security proof, as discussed further in Ref. Section 6.2.2. We highlight however that this result by itself does not constitute a security proof for a QKD protocol. This is firstly because of issues similar to those raised in Section 4.5 — for instance, one needs to address the question of whether the states being analyzed are conditioned on various events (such as the protocol accepting or the key length being chosen to be a particular value); also, it is usually impossible to find a constant upper bound $N_{\mathrm{const}}^{\mathrm{err}}$ on the number of phase errors, similar to the problem of finding a constant lower bound $\kappa_{\mathrm{const}}$ on the smooth min-entropy. Furthermore, since the proof technique only analyzes hashing the string $Z_{A}^{m}$ specifically, to analyze a QKD protocol it must be explicitly justified that the pre-amplification string $\mathbf{S}$ can be written in the form $\mathbf{S}=Z_{A}^{m}$, in some state of the form $\mathcal{E}_{Z}(\rho_{A^{m}B^{m}\mathbf{E}})$ such that the number of phase errors can also be suitably analyzed. Approaches to handle these issues have been presented in [60, 62, 88], though care is needed in interpreting their results; we discuss this further in Section 6.2.1.
@@
    This proof technique requires that the hash family used is linear. Moreover, the proof technique requires that the hash family is dual $2$-universal [89]. These conditions are rather technical and unique to this proof technique, hence we do not discuss them further in detail — we note simply that Toeplitz matrices and random binary matrices are both dual $2$-universal hash families.
@@ 6.2.2 Comparisons with EUR
As mentioned earlier, there are some similarities between the EUR-based security proof approaches and the phase error correction-based security proof approaches, though there are also some differences, which we shall also elaborate on here. First, we note that Ref. [92, Theorem 1] showed that a bound on the smooth min-entropy of a classical string conditioned on some side-information gives a bound on the failure probability of a corresponding phase error correction protocol, in an attempt to bridge the gap between Leftover Hashing Lemma based approaches and phase error correction based approaches. However, this connection is only made in terms of statements similar to the result in Refs. [85, 86], which as mentioned above, does not constitute a full security proof for a QKD protocol (as it is only a statement about some abstract state assumed to satisfy certain properties, rather than the pre-amplification state in a QKD protocol).
@@ 6.3.1 Security proofs against IID collective attacks
    To find such a bound, one can invoke the asymptotic equipartition property (AEP) [44], which lower bounds the smooth min-entropy of an IID state in terms of the von Neumann entropy of a single round, along with some “finite-size correction” terms. This reduces the task to lower bounding the von Neumann entropy produced in a single round from any state $\rho_{ABE}$ constrained to be in the set $S_{\varepsilon_{\mathrm{AT}}}$, which can be addressed using convex optimization techniques [96, 95].
@@
In the above outline, the smooth min-entropy being bounded is that of the state without conditioning on the protocol accepting. [^45] The difficulty with analyzing the state conditioned on accepting is that it may no longer be IID, hence the AEP cannot be applied. To rigorously address the issue mentioned in Section 4.5 regarding how this conditioning affects entropies, one can use [4, Lemma 10] to relate this to the smooth min-entropy of the subnormalized state conditioned on accepting, after which a suitable version of the Leftover Hashing Lemma [4, Proposition 9] can be applied.
@@
For security proofs that are instead based on Rényi entropies, one can use [97, Lemma B.5] to relate the Rényi entropies of the states with and without conditioning on accepting. That lemma is slightly easier to use in that it applies to the normalized conditional state, rather than the subnormalized one; however, the lower bound it gives has an explicit dependence on the accept probability, as briefly noted in Section 4.5.2.
@@
    • Source-replacement scheme: The challenge here was that the source-replacement scheme required Alice to have a fixed marginal state (which was used in the security proof against IID attacks), and needs to be accounted for in the mixture of IID states $\tau_{A^{n}B^{n}}$. This was addressed through the usage of a modified de Finetti theorem [101, Corolary 3.2] which allowed the mixture to be over all IID states that had the same marginal as $\rho_{A^{n}B^{n}}$. [^50] Note that for this to hold, $\rho_{A^{n}}$ must also be IID. This is equivalent to assuming that Alice’s prepares random states in an IID fashion. Also, this work improved on the cost $g_{n,x}$ in the de Finetti theorem stated in Ref. [101].
@@
    For fixed-length protocols, this work showed that if the IID security proof proceeds specifically by bounding the smooth min-entropy of all unfiltered states (satisfying Alice’s reduced-state constraint) as described in Section 4.5.1, then it can be validly lifted to a security proof against coherent attacks, though with larger effects on the security parameter as compared to the initial claim in Ref. [48]. More generally, for either fixed-length or variable-length protocols, this work showed that such a lift is possible even without that condition, but the effect on the security parameter is slightly larger.
In addition to addressing these limitations, this work also derived some improvements to the finite-size terms appearing in the de Finetti theorem, reducing the impact of those terms on the finite-size key rates. Note that although this work describes a lift for decoy-state protocols, it still requires a finite-size, IID security proof for decoy-state protocols. Although we do not believe this to be significantly different from other IID security proofs, we do not know of any currently published self-contained decoy-state security proof.

### v2

@@ 5.4.6 Error verification and the correctness condition
As previously mentioned in Section 4.3, care is necessary when interpreting the guarantees of error verification. In particular, a common misconception is to claim that conditioned on the hashes matching, the probability of the pre-amplification strings differing is small. This is not guaranteed by (for instance) error verification based on $\delta$-AU2 hashing, which only ensures that conditioned on the pre-amplification strings being different, then the probability that their hashes match is small—these two conditional probabilities can have extremely different values, but conflating them is a common mistake.
We now present here an explicit counterexample to the former claim, i.e. we present an attack such that even conditioned on the hashes matching, the pre-amplification strings differ with high probability. The attack is very simple, though showing that it indeed yields the claimed property is slightly elaborate. For simplicity, let us focus on protocols where Alice’s pre-amplification string is a uniformly random string of $m$ bits corresponding exactly to her signal state choices in $m$ generation rounds, and one-way error correction takes place from Alice to Bob with a string of length upper bounded by $\xi m$ for some constant $\xi<1$. Now consider a trivial attack where Eve simply intercepts and discards all states that Alice sends, and forwards some arbitrary unrelated states to Bob.
Observe that with this, all registers that Bob holds before error correction (including basis announcements) are completely independent of Alice’s pre-amplification string, and thus his probability of guessing that string at that point is at most $2^{-m}$. By chain rules for guessing probability [^37] Basically, Eq. (17) with the smoothing parameter $\overline{\varepsilon}$ set to zero, noting that min-entropy has an interpretation in terms of guessing probability [58]., his probability of guessing Alice’s pre-amplification string after receiving the error correction string is then at most $2^{-(1-\xi)m}$. [^38] Here for simplicity we suppose that for each $m$ an error correction string of some fixed length (smaller than $\xi m$) is used; if the string length can vary up to $\xi m$, then we would simply obtain a bound of $2^{-(1-\xi)m+1}$ instead, since there are only at most $2^{\xi m+1}$ such strings. Let $\Omega_{\mathrm{diff}}$ and $\Omega_{\mathrm{EV}}$ respectively denote the events that the pre-amplification strings differ and that error verification accepts, so the probability of interest (e.g. in Eq. (14)) is just $\Pr[\Omega_{\mathrm{diff}}|\Omega_{\mathrm{EV}}]$. With this we observe that for the described attack,
$\Pr[\Omega_{\mathrm{diff}}\text{ AND }\Omega_{\mathrm{EV}}]$ $=\Pr[\Omega_{\mathrm{EV}}|\Omega_{\mathrm{diff}}]\Pr[\Omega_{\mathrm{diff}}]$
$\geq\Pr[\Omega_{\mathrm{EV}}|\Omega_{\mathrm{diff}}](1-2^{-(1-\xi)m}),$ (18)
and (writing $\neg$ to denote negation of an event)
$\Pr[\Omega_{\mathrm{EV}}]$ $=\Pr[\Omega_{\mathrm{EV}}|\Omega_{\mathrm{diff}}]\Pr[\Omega_{\mathrm{diff}}]$
$+\Pr[\Omega_{\mathrm{EV}}|\neg\Omega_{\mathrm{diff}}]\Pr[\neg\Omega_{\mathrm{diff}}]$
$\leq\Pr[\Omega_{\mathrm{EV}}|\Omega_{\mathrm{diff}}]\Pr[\Omega_{\mathrm{diff}}]+(1)(1-\Pr[\Omega_{\mathrm{diff}}])$
$\leq 1+\left(\Pr[\Omega_{\mathrm{EV}}|\Omega_{\mathrm{diff}}]-1\right)(1-2^{-(1-\xi)m})$
$=\Pr[\Omega_{\mathrm{EV}}|\Omega_{\mathrm{diff}}](1-2^{-(1-\xi)m})+2^{-(1-\xi)m},$ (19)
where we used $\Pr[\Omega_{\mathrm{EV}}|\Omega_{\mathrm{diff}}]-1\leq 0$ in the third inequality. This gives
$\Pr[\Omega_{\mathrm{diff}}|\Omega_{\mathrm{EV}}]$ $=\frac{\Pr[\Omega_{\mathrm{diff}}\text{ AND }\Omega_{\mathrm{EV}}]}{\Pr[\Omega_{\mathrm{EV}}]}$ (20)
$\geq\frac{\Pr[\Omega_{\mathrm{EV}}|\Omega_{\mathrm{diff}}](1-2^{-(1-\xi)m})}{\Pr[\Omega_{\mathrm{EV}}|\Omega_{\mathrm{diff}}](1-2^{-(1-\xi)m})+2^{-(1-\xi)m}}$
$=\frac{1-2^{-(1-\xi)m}}{1-2^{-(1-\xi)m}+\frac{2^{-(1-\xi)m}}{\Pr[\Omega_{\mathrm{EV}}|\Omega_{\mathrm{diff}}]}}.$ (21)
For simplicity, suppose furthermore that error verification is performed with a strongly universal2 hash [118], which has the property that $\Pr[\Omega_{\mathrm{EV}}|\Omega_{\mathrm{diff}}]=2^{-\ell_{\mathrm{hash}}}$ where $\ell_{\mathrm{hash}}$ is the hash length. Then the last line in the above bound limits to $1$ as $m\to\infty$ (taking $\ell_{\mathrm{hash}}$ to be constant, or sublinear with respect to $m$). Hence $\Pr[\Omega_{\mathrm{diff}}|\Omega_{\mathrm{EV}}]$ becomes arbitrarily close to $1$ at large $m$, as claimed.
While we have focused on a simple example here, the overall principle should generalize to other cases. For instance, while we considered strongly universal2 hashing as an example, the same conclusion holds for any error verification procedure in which $2^{-(1-\xi)m}$ is “vanishingly small” compared to $\Pr[\Omega_{\mathrm{EV}}|\Omega_{\mathrm{diff}}]$ at large $m$ (more rigorously: their ratio limits to zero as $m\to\infty$). We expect this to indeed be the case for most error verification procedures, including the alternative options mentioned in Section 3.4.2 that are implicitly built into the acceptance test — this is because in order to obtain convergence to the asymptotic key rate, the number of bits announced for the acceptance test or error verification usually has to be sublinear in $m$, and it seems unlikely that $\Pr[\Omega_{\mathrm{EV}}|\Omega_{\mathrm{diff}}]$ can be of comparable magnitude to $2^{-(1-\xi)m}$ in that case. [^39] A possible exception is if error correction is performed with strings of variable length based on some announced data, such that the maximum possible length is not upper bounded by $\xi m$ for any constant $\xi<1$. However, this forces the protocol to be a variable-length protocol, and thus it must be accompanied by a rigorous variable-length security proof.
In summary, it would be incorrect for a security proof to claim at any point that $\Pr[\Omega_{\mathrm{diff}}|\Omega_{\mathrm{EV}}]$ is small (unless the protocol has sufficient additional features to rule out the above counterexample). With this in mind, we re-emphasize however that as observed in Section 4.3, $\Pr[\Omega_{\mathrm{diff}}|\Omega_{\mathrm{EV}}]$ is not in fact relevant when proving that the correctness condition holds. This is because the correctness condition is not phrased in terms of this conditional probability—it only requires that the joint probability of the keys differing and the protocol accepting is small, which can be rigorously bounded using only the guarantees of $\delta$-AU2 hashing, as noted in Eq. 13. The framework of composable security then ensures that we have the desired composable properties of the protocol, despite the fact that $\Pr[\Omega_{\mathrm{diff}}|\Omega_{\mathrm{EV}}]$ can be large.
@@ 5.5 Variable-length protocols
For security proofs of variable-length protocols see Refs. [114, 75] for an approach involving the use of the postselection technique [22], see [24, Supplementary Note A], [113, Appendix B, F] for a security proof in the EUR framework, and [53, Chapter 3] [45] for the phase error correction based methods.
@@ 5.6.1 Resources used in QKD
While the RNGs and the authenticated classical channels can in principle both be constructed with information-theoretic security, for some QKD implementations one might instead wish to use constructions that only have computational security. There exist works considering this possibility and its connection to a concept known as everlasting security (see e.g. Ref. [82, Section VIII.D.4]), but that will not be a focus of this work.
@@ 5.6.2 Variable-length of pre-amplification string
In theoretical analysis, sifting is typically modeled as setting the pre-amplification string register to a 00 or $\bot$ for each round that is discarded, which leads to the pre-amplification string register having a fixed length. Privacy amplification is then implemented by universal2 hashing on this register (of fixed-length) to some $l$ bits. Note that if the discarded rounds are set to $\bot$, then the hashing procedure is no longer applied to binary strings.
@@
In practice, sifting is often implemented by physically discarding the round, which leads to the pre-amplification string register having a variable number of bits. Then, privacy amplification is done by first looking at the number of bits in the pre-amplification string, and then implementing universal2 hashing from those many bits to $l$ bits. This approach is preferred since it allows users to apply hashing on smaller, binary strings.
This difference constitutes a clear gap between the theoretical model and the practical implementation. The gap is further exacerbated by the fact that the implemented protocol does not represent a valid universal2 hashing procedure on the pre-amplification string register [^40] For instance, consider two all-zero strings of different lengths. If we were to look at the number of bits in each string and then apply a universal2 hashing procedure (such as Toeplitz hashing) for that string length to that string, the output would always be all-zero strings in both cases, i.e. the collision probability would be $1$ even though the input strings are different.. This issue was first pointed out and addressed in Ref. [114, Section V], where a variant of the Leftover Hashing Lemma that is applicable to the variable-length pre-amplification strings case was proven. In particular, it was shown that as long as the positions of the discarded rounds are publicly announced, the theoretical analysis of the following three scenarios are equivalent:
@@
Note that the above problem does not arise in EUR and phase error correction based proof techniques. This is because in these frameworks, the pre-amplification string is only generated from detected signals. If the pre-amplification string is of variable length, then one can separately apply the Leftover Hashing Lemma for each possible length, and obtain a security proof similar to the one for variable-length protocols [113, Remark 17]. Alternatively, some protocols choose a random subset of fixed length for the pre-amplification string on which privacy amplification is done.
@@ 6.1 Proofs based on entropic uncertainty relations
Additionally, this work requires the source preparation to be perfect. However, there are several follow up works on the topic of source imperfections, that provide methods to obtain a suitable bound on the phase error rate, even in the presence of source imperfections [25, 125, 42, 99, 23, 96, 97, 123, 79, 77]. It also considers active basis choice by Bob, and assumes that the probability of detection in Bob’s setup is independent of basis. This assumption is a core issue for all EUR and phase error correction (see Section 6.2) based proofs for standard QKD, and was removed recently in Ref. [113].
6.1.2 Ref. [88]
@@ 6.2 Phase error correction approach
Refs. [56, 55] showed that if there exists some constant $N_{\mathrm{const}}^{\mathrm{err}}\in\mathbb{N}$ that bounds the number of phase errors with high probability, then hashing [^43] There are some technical points regarding the hash families considered in this technique; we shall discuss this point shortly. the string $Z_{A}^{m}$ in the state $\rho_{Z_{A}^{m}\mathbf{E}}^{(Z)}$ to a length determined by $N_{\mathrm{const}}^{\mathrm{err}}$ results in the output being nearly uniform and decoupled from $\mathbf{E}$, roughly analogous to the Leftover Hashing Lemma. [^44] There are a number of similarities between this approach and EUR-based approaches when constructing a QKD security proof, as discussed further in Sec. Section 6.2.2. We highlight however that this result by itself does not constitute a security proof for a QKD protocol. This is firstly because of issues similar to those raised in Section 4.5 — for instance, one needs to address the question of whether the states being analyzed are conditioned on various events (such as the protocol accepting or the key length being chosen to be a particular value); also, it is usually impossible to find a constant upper bound $N_{\mathrm{const}}^{\mathrm{err}}$ on the number of phase errors, similar to the problem of finding a constant lower bound $\kappa_{\mathrm{const}}$ on the smooth min-entropy. Furthermore, since the proof technique only analyzes hashing the string $Z_{A}^{m}$ specifically, to analyze a QKD protocol it must be explicitly justified that the pre-amplification string $\mathbf{S}$ can be written in the form $\mathbf{S}=Z_{A}^{m}$, in some state of the form $\mathcal{E}_{Z}(\rho_{A^{m}B^{m}\mathbf{E}})$ such that the number of phase errors can also be suitably analyzed. Approaches to handle these issues have been presented in [45, 57, 44], though care is needed in interpreting their results; we discuss this further in Section 6.2.1.
@@
    Refs. [56, 55] only studied a specific hashing procedure; however, subsequent work [110] generalized the result to a broader class of hashing procedures known as almost-dual-universal2 hashing — we do not discuss the technical details here, but simply note that this includes universal2 hashing as a special case. It does require that the hash family consists of linear functions, though we note this is indeed satisfied by commonly used procedures such as Toeplitz hashing.
@@ 6.2.2 Comparisons with EUR
As mentioned earlier, there are some similarities between the EUR-based security proof approaches and the phase error correction-based security proof approaches, though there are also some differences, which we shall also elaborate on here. First, we note that Ref. [109, Theorem 1] showed that a bound on the smooth min-entropy of a classical string conditioned on some side-information gives a bound on the failure probability of a corresponding phase error correction protocol, in an attempt to bridge the gap between Leftover Hashing Lemma based approaches and phase error correction based approaches. However, this connection is only made in terms of statements similar to the result in Refs. [56, 55], which as mentioned above, does not constitute a full security proof for a QKD protocol (as it is only a statement about some abstract state assumed to satisfy certain properties, rather than the pre-amplification state in a QKD protocol).
@@ 6.3.1 Security proofs against IID collective attacks
    To find such a bound, one can invoke the asymptotic equipartition property (AEP) [104], which lower bounds the smooth min-entropy of an IID state in terms of the von Neumann entropy of a single round, along with some “finite-size correction” terms. This reduces the task to lower bounding the von Neumann entropy produced in a single round from any state $\rho_{ABE}$ constrained to be in the set $S_{\varepsilon_{\mathrm{AT}}}$, which can be addressed using convex optimization techniques [120, 38, 62].
@@
In the above outline, the smooth min-entropy being bounded is that of the state without conditioning on the protocol accepting. [^50] The difficulty with analyzing the state conditioned on accepting is that it may no longer be IID, hence the AEP cannot be applied. To rigorously address the issue mentioned in Section 4.5 regarding how this conditioning affects entropies, one can use [105, Lemma 10] to relate this to the smooth min-entropy of the subnormalized state conditioned on accepting, after which a suitable version of the Leftover Hashing Lemma [105, Proposition 9] can be applied.
@@
For security proofs that are instead based on Rényi entropies, one can use [31, Lemma B.5] to relate the Rényi entropies of the states with and without conditioning on accepting. That lemma is slightly easier to use in that it applies to the normalized conditional state, rather than the subnormalized one; however, the lower bound it gives has an explicit dependence on the accept probability, as briefly noted in Section 4.5.2.
@@
    • Source-replacement scheme: The challenge here was that the source-replacement scheme required Alice to have a fixed marginal state (which was used in the security proof against IID attacks), and needs to be accounted for in the mixture of IID states $\tau_{A^{n}B^{n}}$. This was addressed through the usage of a modified de Finetti theorem [33, Corollary 3.2] which allowed the mixture to be over all IID states that had the same marginal as $\rho_{A^{n}B^{n}}$. [^55] Note that for this to hold, $\rho_{A^{n}}$ must also be IID. This is equivalent to assuming that Alice’s prepares random states in an IID fashion. Also, this work improved on the cost $g_{n,x}$ in the de Finetti theorem stated in Ref. [33].
@@
    For fixed-length protocols, this work showed that if the IID security proof proceeds specifically by bounding the smooth min-entropy of all unfiltered states (satisfying Alice’s reduced-state constraint) as described in Section 4.5.1, then it can be validly lifted to a security proof against coherent attacks, though with larger effects on the security parameter as compared to the initial claim in Ref. [22]. More generally, for either fixed-length or variable-length protocols, this work showed that such a lift is possible even without that condition [75, Section V. A.], but the effect on the security parameter is slightly larger.
In addition to addressing these limitations, this work also derived some improvements to the finite-size terms appearing in the de Finetti theorem, reducing the impact of those terms on the finite-size key rates. Note that although this work describes a lift for decoy-state protocols, it still requires a finite-size, IID security proof for decoy-state protocols.
6.3.4 Ref. [51]
Recently, Ref. [51] presented an IID security proof for generic decoy-state protocols, together with a lift to coherent attacks using the postselection technique. The two significant contributions of this work, besides the IID analysis for decoy-state protocols and the explicit key rate computations against coherent attacks from this approach, are a tightened statistical analysis and reduced finite-size corrections. The tightened statistical analysis improves key rates, and is especially relevant in the case of very small security parameters, which are required in the postselection technique. Furthermore, the reduced second-order finite-size term in the key length scales with the square root of number of detected events, rather than the total number of signals, leading to improved finite-size performance.